Packets of milk powder produced by a machine were found to have a normal distribution with a mean mass of 650g and a standard deviation of 10g.

Find the probability that a packet selected at random will have a mass between 620g and 655g
If 500 packets are selected at random, how many of them will have a mass of more than 660g?
It is found that 10% of packets of milk powder will have a mass of less than k Calculate k.
1
Expert's answer
2019-08-21T03:59:18-0400

Suppose "X" is a variable representing the mass of the packet of milk powder. "X" is normally distributed with mean of 650 g and standard deviation of 10 g: 


"X\sim N(650, 10^2)"

If "X\sim N(\mu, \sigma^2)", then "Z={X-\mu \over \sigma}\sim N(1, 0)"

The probability that a packet selected at random will have a mass between 620g and 655g 


"P(620<X<655)=P({620-650 \over 10}<Z<{655-650 \over 10})=""=P(-3<Z<0.5)=F(0.5)-F(-3)=""=F(0.5)-(1-F(3))=0.691462-0.001350=""=0.690112"

If 500 packets are selected at random, how many of them will have a mass of more than 660g? 

Given that "N=500."

"P(X>660)=P(Z>{660-650 \over 10})=1-F(1)=""=1-0.841345=0.158655""n=N\cdot P(X>660)=500\cdot 0.158655=79"

It is found that 10% of packets of milk powder will have a mass of less than k .


"P(X<k)=0.1""P(X<k)=P(Z<{k-650 \over 10})=0.1""{k-650 \over 10}=-1.281552"

"k=650-10\cdot1.281552\approx637"

The probability that a packet selected at random will have a mass between 620g and 655g  is "0.690112."

If 500 packets are selected at random, then 79 of then will have a mass of more than 660g.

It is found that 10% of packets of milk powder will have a mass of less than k=637 g .